# Joint Diffusion models in Continual Learning

### v1

@@ 5.3 Additional Experiments
Table 3: Ablation study depicting the importance of individual components. Our method greatly relies on the synergy of joint modeling with knowledge distillation.
@@
In this work, we propose a new generative replay method for continual learning. Apart from the main idea of our approach which is joint diffusion and classification modeling, we combine it with two important techniques: knowledge distillation and two-stage training. In this section, we ablate the impact of each component. In table Tab. 3 we show the results for the model trained without individual components. As visible, the performance of our method is attributed mostly to the synergy of the proposed joint modeling with a knowledge distillation technique.
Stability of the JDCL’s representation space in an incremental setup
Figure 4: Accuracy of the kNN-clasifier trained on representations provided by the model after the first task of class-incremental CIFAR100 setup with 5 tasks
One of the most important features of our methods is its ability to learn a meaningful representation space shared by the generative and discriminative parts of the model. To better understand how this representation space changes during the class-incremental training we propose to measure the drift of representations, with the following experiment on a 5 task CIFAR100 setup. We first extract data representations from two parts of our UNet model, the H-space inside of the generative part of our model, and the penultimate layer of the classifier head. We train a kNN-classifier on top of those representations provided by the model after the first task. Then, w evaluate how the accuracy of this classifier changes when the representations for data from the first task are extracted with models trained incrementally on later tasks. We present the results in Fig. 4. We observe that the representations extracted from the UNet change significantly slower than those from the classifier. This means, that representations that are not used for classification drift significantly slower than those that are classifier-specific,what highlights the benefits of joint modeling in a continual setup.
@@
Figure 5: Accuracy of a logistic regressor trained on representations learned with different self-supervised methods on class-incremental CIFAR100. Results of competing approaches from [15]
Finally, in this section, we show that our method is able to learn meaningful data representation in a continual setup. To that end, we follow the evaluation approach common in self-supervised continual learning techniques, as presented in [15]. In particular, we report the accuracy of a logistic regressor trained on the features provided by our model i.e. on the set of representations of the UNet model, pooled into a single vector. In Fig. 5 we compare our model with self-supervised continual methods. We can observe that the representations learned by our model are indeed very informative. Moreover, we find that JDCL outperforms other approaches on the CIFAR100 setup with 5 tasks. However, it is important to note that the other methods are self-supervised and thus, did not use labels during training.
@@ 6 Conclusions
In this work, we introduce JDCL – a new continual learning method that employs joint-diffusion modeling for generative replay. We propose a shared parametrization that mitigates the problem of knowledge transfer in generative replay. In our experimental section, we show that JDCL outperforms recent generative replay techniques in standard class incremental scenarios. Moreover, we propose the adaptation of our method to semi-supervised learning where we are able to outperform buffer-based approaches. Finally, we show additional experiments that the performance of our method should be attributed to its ability to learn and maintain useful representations.
Appendix
Appendix A Implementation details
We train diffusion using linear noise schedulers over 1000 steps and use DDIM with 250 steps to generate replay samples. In particular, we generate 1000 and 400 samples per class in CIFAR10 and CIFAR100 respectively. For CIFAR10 setups, we train our local model for 60000 steps, and for CIFAR100 setups, we train it for 100000 steps. For the second task, we train the global model for 60000 and 100000 steps for CIFAR10 and CIFAR100 respectively. For all the following tasks, we train for 30000 steps for CIFAR10 and 70000 steps for CIFAR100. Additionally, following [10], for the first 10000 steps of every global model training, we do not calculate the classification loss and optimize only the generative objective. We use a single AdamW optimizer with a learning rate set to 0.0002 and a batch size of 256. For all experiments, we set the classification loss scales $\alpha$ and $\alpha_{KD}$ to 0.001, and the $\beta$ hyperparameter that scales joint-diffusion loss to $0.01$.
We define three sets of data augmentations. The first one consists of flipping, resized cropping, and normalization to range $[-1,1]$, and is used to train the denoising model. The second, strong augmentation set includes flipping, cropping, the RandAugment [8] augmentation policy, and normalization to range $[-1,1]$. The third and final, weak augmentation set consists of flipping, cropping, and normalization to range $[-1,1]$. We use the strong augmentation set to perturb the images for the classifier in the supervised setups.
Additionally, following [60], we use our strong and weak augmentation sets to compute the semi-supervised loss. We also set the $\tau$ threshold to $0.95$ and we set the supervised-to-unsupervised batch size ratio to 7, which corresponds to a supervised batch size of 32 and an unsupervised batch size of 224.
Appendix B Representation space in a semi-supervised setup
In this work, we introduce a new method of training joint diffusion model in a semi-supervised setup. In this section, we show that this approach allows our model to learn shared representations that are meaningful for both generative and discriminative objectives. To that end, we perform a visual comparison of the representation spaces in Fig. 6. We compare our method with a state-of-the-art semi-supervised method FixMatch [60], and a baseline classifier model that we train only on the labeled data. To extract the representations we average pool the UNet representations into a single vector in our approach, while for the other two models, we take the output of the second to last layer. We observe that our approach is able to group the labeled and unlabeled data from the same class into similar regions of the embedding space shared between the classifier and diffusion model, similar to the state-of-the-art FixMatch approach.
Figure 6: Visualization of the representation spaces (umap) of different methods trained on a partially labeled dataset. Labeled data samples are marked with an "X", while unlabeled data points are marked with circles. All the models were trained on the first 10 classes of CIFAR100 with 5% of the data being labeled.
Appendix C Limitations and Broader impact
Our solution is based on the generative replay continual learning strategy that is known to suffer from the big limitation related to the need for training and sampling from the generative model. Therefore, our solution is by definition computationally more expensive than competing approaches that rely on the buffer of examples. However, compared to other generative replay approaches, we decrease our computational cost by training a single joint model instead of a separate classifier.
This work tackles the problem of continual learning, which might have some potential societal impact, mostly generic for machine learning research. However, our approach is based on diffusion-based generative models that have some known and many hidden limitations that might lead to biases and potential security risks.
Appendix D Computational resources
To conduct our experiments we used NVIDIA A100-SXM4-40GB GPU and an AMD EPYC 7742 64-Core CPU. For every run, we utilized exactly one GPU. Each experiment in a continual CIFAR10 setup with 5 tasks took around 100 hours, experiment on CIFAR100 with 5 tasks needed around 200 hours, and the experiment on CIFAR100 with 10 tasks needed approximately 400 hours. To run all experiments disclosed in the paper we have used around 5000 computational hours. In total, the whole research required around 12,000 hours of computations.

### v2

@@ 5.3 Additional Experiments
In this work, we propose a new generative replay method for continual learning. Apart from the main idea of our approach which is joint diffusion and classification modeling, we combine it with two important techniques: knowledge distillation and two-stage training. In this section, we ablate the impact of each component. Notably, for experiments without joint modeling, we use a separate diffusion model with the same architecture as the joint diffusion model, and ResNet-18 as a classifier. We train and distill them in the same manner as the joint model. In table Tab. 3 we show the results for the model trained without individual components. As visible, the performance of our method is attributed to the synergy of the proposed joint modeling, knowledge distillation technique, and two-stage training.
Table 3: Ablation study on CIFAR10/5 depicting the importance of individual components. Our method greatly relies on the synergy of joint modeling with knowledge distillation.
@@
In this section, we show that our method is able to learn meaningful data representation in a continual setup. To that end, we follow the evaluation approach common in self-supervised continual learning techniques, as presented in [15]. In particular, we report the accuracy of a logistic regressor trained on the features provided by our model i.e. on the set of representations of the UNet model, pooled into a single vector. In Fig. 4 we compare our model with self-supervised continual methods. We can observe that the representations learned by our model are indeed very informative. Moreover, we find that JDCL outperforms other approaches on the CIFAR100 setup with 5 tasks. However, it is important to note that the other methods are self-supervised and thus, did not use labels during training.
Figure 4: Accuracy of a logistic regressor trained on representations learned with different self-supervised methods on CIFAR100. Results of competing approaches from [15]
Table 4: Comparison of the generation quality of JDCL and Deep Generative Replay with standard DDPM after the last task.
Method CIFAR-10/5 CIFAR-100/5
FID Precision Recall FID Precision Recall
DGR 49.56 0.54 0.34 50.36 0.48 0.30
JDCL 32.12 0.60 0.43 39.93 0.71 0.27
Incremental generative setup comparison
Generative-replay-based continual learning methods rely heavily on the underlying generative model to preserve knowledge about the previous tasks. The quality of the samples generated by the method is therefore of utmost importance. In this section, we perform a comparison of JDCL and the diffusion model trained continually on CIFAR10 and CIFAR100, both split into 5 tasks. We present the achieved FID, Precision, and Recall after the last task in Table 4. JDCL outperforms the baseline method on all but one metric. We can observe significant improvement in the precision of generations, which results in generated rehearsal samples being clearly assigned to one of the previous classes bringing less noisy signals to the model. As argued in [18], this trait can be the reason for the improved performance of the continually trained classifier.
Runtime analysis of JDCL for CIFAR-10/5
One of the biggest drawbacks of generative-replay-based techniques is an excessive computational cost associated with the training of the generative model and additional time spent on sampling of the rehearsal examples. However, thanks to the joint modeling our method is very data-efficient. For example, in the case of the CIFAR dataset, at the beginning of each task, we sample only 400 examples per class for CIFAR-100 or 1000 examples per class for CIFAR10. This occupies approximately 13% of the total GPU time spent on each task. The combination of the joint model training additionally allows us to achieve state-of-the-art results with computational costs comparable to other diffusion-based solutions. Moreover, with simple tricks such as training with reduced float precision (JDCL fast) our method achieves performance significantly higher than other approaches, while also having the lowest computational cost. In Table 5, we present a runtime analysis to compare the training times of JDCL with diffusion-based baselines. All methods were trained on the same machine with a single NVIDIA A100 GPU.
Table 5: Runtime analysis of all baseline methods on the CIFAR-10/5 benchmark. Our approach is computationally more efficient than related diffusion-based methods.
Method Time [GPU-hours] Accuracy
DDGR $41.29$ 43.69
DGR diffusion 28.02 59.00
GUIDE 30.60 64.47
JDCL (fast) 27.21 78.10
JDCL 53.83 83.69
@@ 6 Conclusions
In this work, we introduce JDCL – a new continual learning method that employs joint-diffusion modeling for generative replay. We propose a shared parametrization that mitigates the problem of knowledge transfer in generative replay. In our experimental section, we show that JDCL outperforms recent generative replay techniques in standard class incremental scenarios. Moreover, we propose the adaptation of our method to semi-supervised learning where we are able to outperform buffer-based approaches. Finally, we show in additional experiments that the performance of our method should be attributed to its ability to learn and maintain useful representations.